OpenStudy (christos):

1 - 3 + 5 - 7 + 9 - 11 = Find the summation

hartnn (hartnn):

treat them as 2 summations, (1+5+9)- (3+7+11) 1+5+9 3+7+11 then subtract...

OpenStudy (christos):

= 26 ?

hartnn (hartnn):

no.... 1+5+9=15 3+7+11=21 15-21=-6

hartnn (hartnn):

and there are only 6 terms?? you could have use calculator or calculated it mentally.

hartnn (hartnn):

alternatively, you could have taken 2 terms at a time, 1-3 =-2, 5-7=-2, 9-11=-2 and -2 + -2+ -2 = -2*3 =-6

OpenStudy (skullpatrol):

Do you know the definition of subtraction as being adding the opposite?

OpenStudy (skullpatrol):

@Christos ?

OpenStudy (skullpatrol):

The simplest way is to use the order of operations and do the additions and subtractions in order from left to right.

OpenStudy (skullpatrol):